following equation defines a relation (x,y). Write the domain and range of each relation.
x^2/9 + y^2/16 = 1

Answer by stanbon(48581)

Write the domain and range of each relation.
x^2/9 + y^2/16 = 1
(y^2/16) = 1 - (x^2/9)
y^2 = 16[1-(x^2/9)]
y = +-4sqrt[1-(x^2/9)]
-----------------------
Domain: All Real Numbers between -3 and 3
Range: All Real Numbers between -4 and 4
==========================================